OpenStudy (anonymous):

solve this ststem of equations by using the elimination method: 13x-3y=-50 12x+5y=16

OpenStudy (anonymous):

please show work

OpenStudy (amistre64):

how do we start elinating stuff?

OpenStudy (amistre64):

*eliminating stuff, that is

OpenStudy (amistre64):

if i have $5 and want to get rid of it, what should I do?

OpenStudy (anonymous):

give it to me?

OpenStudy (amistre64):

exactly, I should perform an operation whereby I get rid of it :)

OpenStudy (amistre64):

we are adding 2 equations, and we want to add them up so that we get rid of one of the variables so that we aint got to mess with it

OpenStudy (anonymous):

please just answer the question, u are getting me confused

OpenStudy (amistre64):

this is the answering part ....

OpenStudy (wasiqss):

listen elimination i can do pretty easily bt the question will go into decimal and hard to solve so betta do by substitution

OpenStudy (anonymous):

I like elimination more; less steps

OpenStudy (amistre64):

this is the means whereby we encourage you to study up on how the process is performed in order that you may gain a better understanding of the principles involved.

OpenStudy (anonymous):

i know how to do substitution, but i want to learn how to do it in elimination

OpenStudy (amistre64):

i like cramer myself, it is the culmination of elimination and soo much easier

OpenStudy (amistre64):

would you agree that we can multiply or divide an equation and not change its inherent properties?

OpenStudy (anonymous):

?

OpenStudy (wasiqss):

ok multiply the second second eq with .8

OpenStudy (wasiqss):

then y of both equation will b equal hence they will cancel out

OpenStudy (espex):

@amistre64 Evenly to both sides, yes. :)

hero (hero):

wasiqss, the equations don't necessarily have to be multiplied by decimals

OpenStudy (amistre64):

of course ;)

OpenStudy (wasiqss):

it wud get much higher num then, so easiest is multiply by .8

OpenStudy (amistre64):

the foundation to elimination is knowing that scaling an equation doenst change it properties: 3x=9 is the same as 5(3x=9): or simply 15x=45

OpenStudy (wasiqss):

9.6x+3y=1.28 13x-3y=-50

OpenStudy (anonymous):

65x - 15y = -250 36x + 15y = 48 101x = -202 x = -2 y = 8

OpenStudy (wasiqss):

gud one money bird

OpenStudy (anonymous):

thankyou moneybird

hero (hero):

I personally hate the elimination method. Setting the equations equal to each other by writing both in terms of y is better.

hero (hero):

Morning rld

OpenStudy (anonymous):

Hey